OpenStudy (anonymous):

There were 190 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $530. How many Juniors bought tickets?

OpenStudy (anonymous):

13. you need two equations here so 2 dollars for seniors = 2s and 3.5 dollars for juniors = 3.5j so your first equation would look like- 2s + 3.5j = 530 then seniors and juniors combined is going to total 190 so your second equation would look like- s + j = 190. In order to complete this equation you have to write s in terms of j or vice versa so because the second problem is simpler I am going to use it s = 190 - j so 2(190-j) + 3.5j = 530 when you simplify this equation you find that j = 100. Therefore there were 100 juniors at the dance.